OpenStudy (anonymous):

Describe each of the following properties of the graph of the Cosine Function, f(theta) = cos(theta) and relate the property to the unit circle definition of cosine. Amplitude Period Domain Range x-intercepts I don't exactly want the answers I want to know how I would the out the answer, and I notes but I don't understand them

OpenStudy (tkhunny):

Standard, unaltered, just write down the properties. It is a memorization project... \(f(x) = \cos(\theta)\) Amplitude: 1 Period: \(2\pi\) Domain: \(\mathbb{R}\) (That's the Real Numbers) Range: [-1,1] x-intercepts: \(\pi/2 + k\pi\;for\;k\in\mathbb{Z}\) (That's the Integers.) Get this stuck in your head. You will need this basic information to solve the much more complicated ones to come.

OpenStudy (anonymous):

@tkhunny thank you

OpenStudy (tkhunny):

Suomalainen?